# CEM: A Data-Efficient Method for Large Language Models to Continue Evolving From Mistakes

### v6

@@ Abstract
As world knowledge advances and new task schemas emerge, Continual Learning (CL) becomes essential for keeping Large Language Models (LLMs) current and addressing their shortcomings. This process typically involves continual instruction tuning (CIT) and continual pre-training (CPT) to enable these models to adapt to novel tasks and acquire critical knowledge. However, collecting sufficient CPT data and efficiently bridging knowledge gaps remain significant challenges. Inspired by the ‘summarizing mistakes’ strategy, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient approach aiming to collect CPT data and continually improve LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge. To further optimize data usage and mitigate forgetting, we introduce a novel training paradigm that combines CIT and CPT. Experiments show that CEM substantially enhances multiple models’ performance on both in-domain and out-of-domain QA tasks, achieving gains of up to 29.63%. Code and datasets are available on GitHub  [^1] https://anonymous.4open.science/r/cem-BB25.
@@
Haokun Zhao1, Haixia Han2, Jie Shi1, Chengyu Du 1, Jiaqing Liang3 thanks: Corresponding Author, Yanghua Xiao 1,2 Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 {hkzhao23,jshi22}@m.fudan.edu.cn, {haixiahan03,cydu2024}@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
@@ 1 Introduction
With the exponential growth of data and model sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the continual emergence of new knowledge and evolving task requirements necessitate ongoing updates and task-specific adaptations for LLMs. Without these, models risk inducing hallucinations due to knowledge boundaries and may cause task misalignment Huang et al. (2023a). Additionally, addressing and correcting the shortcomings and errors exposed during practical use is crucial. Therefore, Continual Instruction Tuning (CIT) and Continual Pre-training (CPT) are proposed as the primary methods of Continual Learning (CL) to align LLMs with evolving knowledge and tasks, and to improve their shortcomings Wu et al. (2024).
@@
Despite CIT’s smaller data size requirement and lower risk of overfitting compared to CPT, it is limited in new knowledge injection, as Zhou et al. (2023) highlight that instruction tuning is a superficial alignment focused on interaction styles (i.e., task schema) and does not effectively help the model acquire new knowledge. Thus, it cannot address all the issues shown in Figure 1, where LLMs’ poor performance may stem from a lack of requisite knowledge rather than unfamiliarity with the task schema Ren et al. (2023); Zhang et al. (2023b). In contrast, CPT is more effective in addressing the lack of knowledge.
@@
However, using CPT to supplement the model with lacking knowledge still presents several challenges. First, collecting CPT data is challenging as identifying the lacking knowledge a priori while ensuring a sufficiently large data size is difficult. Blind data collection and insufficient data volume may introduce irrelevant or redundant information, weaken generalization, and increase overfitting risks Zhang et al. (2024); Gekhman et al. (2024b). Second, efficiently utilizing CPT data and mitigating degradation are difficult due to weak constraints on parameter updates, leading to larger parameter shifts and a data size gap compared to the initial pre-training.
@@
To address these issues, the most common data collection methods involve using the latest Wikipedia snapshots or new academic papers in specialized fields Cossu et al. (2022); Jin et al. (2022). However, these methods lack focus and efficiency and fail to dynamically update model deficiencies. Additionally, recent work has explored using highly relevant QA pairs for CIT before CPT Jiang et al. (2024) or converting raw CPT corpus into reading comprehension instructions (i.e., CIT data) Cheng et al. (2024b) to mitigate degradation and enhance the efficiency of knowledge absorption from CPT data. Nevertheless, reliance on oracle LLMs and the complexity of mining patterns limit their practicality.
@@
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide (1) a practical pipeline for collecting CPT data based on model mistakes and (2) a novel training set construction paradigm to efficiently utilize these data, enabling continual model improvement while mitigating degradation. To enhance the targetedness and efficiency of data collection, we draw inspiration from the human learning skill of ‘summarizing mistakes’. We argue that a model’s mistakes in ‘exams’ reflect its inherent knowledge deficiencies. Therefore, we collect corpora from the internet that contain background knowledge related to these mistakes to directly address these deficiencies. To efficiently utilize these CPT data and mitigate forgetting, we design a novel training set construction paradigm that integrates parallel CIT and CPT data. In addition, by integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al. (2022), we enable iterative and continual model evolution. The CEM method consistently demonstrates significant improvements across multiple models in extensive experiments, achieving up to a 29.63% accuracy enhancement.
@@ 2 Related Work
Continual pre-training enhances domain knowledge through two approaches: (1) domain-incremental pre-training, which accumulates knowledge across multiple domains, and (2) domain-specific continual learning, which refines a general model into a domain expert by training on domain-specific datasets and tasks. In domain-incremental pre-training, Cossu et al. (2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various downstream tasks. Ke et al. (2023) introduced a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge. Notably, there has been research on continual pre-training for specific domains such as finance Xie et al. (2023) and e-commerce Ma et al. (2023).
@@
Task-incremental Continual Instruction Tuning (CIT) aims to continually fine-tune LLMs on new task-specific instructions. While a straightforward approach involves generating new instruction-tuning data and fine-tuning LLMs directly Wang et al. (2023), this method can lead to catastrophic forgetting of previously learned knowledge and skills Kotha et al. (2024).
@@ 3.1 CEM Method
(1) Search Engine: Search engines are known for high relevance matching and diverse content. We search for extracted knowledge points using Bing and select the top-k results based on relevance, scraping the main textual content as a supplemental corpus. To prevent LLMs from simply memorizing answers, we filter out results that contain both the original question and its answer. Moreover, we use newspaper3k [^3] https://github.com/codelucas/newspaper and readability [^4] https://github.com/buriy/python-readability to exclude irrelevant and spam content.
@@
It should be noted that supplemental training should not be endless, as it may lead to catastrophic forgetting and diminishing returns in accuracy improvement. Subsequent experiments explore the impact of the number of training rounds on model performance.
@@ 4.2 Datasets
We conduct experiments on two question-answering datasets, Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024a). To evaluate CEM’s out-of-domain generalization, we apply CEM on the in-domain dataset Xiezhi and use CMMLU and C-Eval Huang et al. (2023b) as out-of-domain benchmarks.
@@ 4.4 Metrics
ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance. It is calculated by $ER=\frac{R_{k}^{*}-R_{k}^{0}}{R_{k}^{0}},$ where $k$ denotes the task using CEM, $R^{0}_{k}$ is the initial accuracy, and $R^{*}_{k}$ is the accuracy after CEM.
AFR: The Average Forgetting Rate Wang et al. (2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance. It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of tasks $i$ after applying CEM on task $k$.
@@ 5.1 Main Results
Firstly, CEM leads to significant accuracy improvements across all models and tasks. All variations of CEM substantially enhance model performance in both the in-domain Xiezhi task and OOD tasks. For instance, the Llama3-8B-Instruct model achieves accuracy gains of up to 29.63% on the Xiezhi task, 10.92% on the CMMLU task, and 5.40% on the C-Eval task when using CEM-R. This highlights the effectiveness of the CEM method in addressing the model’s knowledge gaps and its capacity to generalize this knowledge to unseen QA tasks.
@@ 5.2 Data Source Analysis
We analyze the impact of the Supplementary Corpus from different sources—Wikipedia, Bing, and their concatenation—on CEM’s performance. To further validate CEM’s generalizability, CMMLU is included as an in-domain task scenario with data proportions consistent with the main experiment.
@@
We find that the CEM method shows robust compatibility across different data sources, consistently improving model performance. However, integrating multiple sources does not necessarily outperform using a single source. As shown in Figure 3, no significant differences in CEM effectiveness emerge when employing Wikipedia, Bing, or their combination. Although the Mix dataset is larger, it often yields lower accuracy than Wiki or Bing, potentially due to data redundancies. Moreover, an excessively large supplemental corpus can dilute the influence of CIT data, reducing its capacity to mitigate catastrophic forgetting, support knowledge extraction, and prevent overfitting.
@@ 5.3 Mechanistic Analysis of Extractive and Review Instruction Components
To clarify the underlying mechanisms by which Extractive Instruction and Review Instruction affect model performance, we conducted experiments on CuteGPT-13B-ift and examined the R2W and W2R metrics. Additionally, we explored the effects of varying $\alpha$ (with values $[0,0.5,1]$) on CEM-R to determine whether more diverse and complex Review Instructions yield improved results. The corresponding results are illustrated in Figure 4 and Figure 5.
We summarize the key observations and insights derived from these analyses:
(1) CEM-E and CEM-R outperform their ablation groups with General Instruction, demonstrating the effectiveness and positive impact of Extractive and Review Instruction on model performance. As shown in Figure 5, the Extractive Instruction enhances W2R performance: compared to CEM-P, CEM-E achieves improvements of 1.45%, 3.22%, and 1.26% across the three models, indicating more effective error correction. In addition, the Review Instruction significantly reduces the R2W metric, reflecting a decreased tendency for the model to forget previously acquired correct knowledge.
@@
(1) CEM-E and CEM-R outperform their ablation groups with General Instruction, demonstrating the effectiveness and positive impact of Extractive and Review Instruction on model performance. Figure 5 shows that Extractive Instruction improves W2R, with CEM-E achieving gains of 1.45%, 3.22%, and 1.26% across the three models compared to CEM-P, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data. This improvement leads to more effective error correction. Additionally, the Review Instruction significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
@@
[…]
Figure 5: The figure presents the R2W and W2R metrics of the models after CEM supplemental training.
@@ 5.4 Multiple Iterations Analysis
To evaluate the iterability of the CEM method and its potential for catastrophic forgetting during multiple training rounds, we conduct three rounds of CEM training on CuteGPT-13B-ift for the Xiezhi task, using consistent training strategies and non-repetitive CIT data of equivalent size in each round. Specifically, the CEM method is applied to the Xiezhi task, while HotpotQA and GSM8K are used to assess reasoning and mathematical skills. The ER metric evaluates performance improvements on Xiezhi, and the AFR metric monitors performance degradation in other tasks. Initially, 2,000 instances from each dataset are sampled for pre-finetuning to establish a baseline. To address forgetting, we implement a random replay strategy, replacing General Instructions with 500 randomly replayed instances from HotpotQA and GSM8K in each training round.
@@
Figure 6 shows that even after three rounds of CEM, performance on Xiezhi improves significantly and maintains an upward trend. The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 75.0% in most cases, highlighting its iterative potential. However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking required by random replay), but more rounds lead to greater forgetting, limiting further iteration ,as seen when CEM-R reaches an ER of 62.4% in round three without replay. CEM-R, with its larger data capacity, achieves significant performance improvements but shows the greatest decline in other tasks, likely due to the increased volume of CPT data diluting the replay effect. Combining CEM with other advanced replay-based methods warrants further investigation.
@@ 6 Conclusion
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes. CEM employs a pipeline that leverages model mistakes to gather extensive CPT data, combined with a novel parallel training set construction paradigm integrating both CIT and CPT data for continuous training. Experiments on various open-source LLMs show that CEM significantly improves answer accuracy and demonstrates strong potential for sustained evolution through multiple iterations.

### v7

@@ Abstract
As world knowledge advances and new task schemas emerge, Continual Learning (CL) becomes essential for keeping Large Language Models (LLMs) current and addressing their shortcomings. This process typically involves continual instruction tuning (CIT) and continual pre-training (CPT) to enable these models to adapt to novel tasks and acquire supplementary knowledge. However, collecting sufficient CPT data and efficiently bridging knowledge gaps remain significant challenges. Inspired by the ‘summarizing mistakes’ strategy, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient approach aiming to collect CPT data and continually improve LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge. To further optimize data usage and mitigate forgetting, we introduce a novel training paradigm that combines CIT and CPT. Experiments show that CEM substantially enhances multiple models’ performance on both in-domain and out-of-domain QA tasks, achieving gains of up to 29.63%. Code and datasets are available on GitHub  [^1] https://anonymous.4open.science/r/cem-BB25.
@@
Haokun Zhao1, Jinyi Han2, Jie Shi1, Chengyu Du 1, Jiaqing Liang3 thanks: Corresponding Author, Yanghua Xiao 1,2 Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University1 Shanghai Institute of Artificial Intelligence for Education, East China Normal University2 Shanghai Key Laboratory of Data Science, School of Data Science, Fudan University3 {hkzhao23,jshi22}@m.fudan.edu.cn, {haixiahan03,cydu2024}@gmail.com, {liangjiaqing,shawyh}@fudan.edu.cn
@@ 1 Introduction
With the exponential growth of data and model sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the continual emergence of new knowledge and evolving task requirements necessitate ongoing updates and task-specific adaptations for LLMs. Without these, models risk hallucinating due to their limited knowledge boundaries and may cause task misalignment Huang et al. (2023a). Additionally, addressing and correcting the shortcomings and errors exposed during practical use is crucial. Therefore, Continual Instruction Tuning (CIT) and Continual Pre-training (CPT) are proposed as the primary methods of Continual Learning (CL) to align LLMs with evolving knowledge and tasks, and to improve their shortcomings Wu et al. (2024).
@@
Although CIT requires less data and poses a lower overfitting risk than CPT, it is less effective at injecting new knowledge. As Zhou et al. (2023) highlight, instruction tuning is a superficial alignment focused on interaction styles (i.e., task schema) and is not well-suited for large-scale knowledge supplementation. This is because scaling up the necessary task-specific instruction data significantly increases complexity in terms of data selection, careful prompt design, and ensuring coverage across extensive knowledge domains. Consequently, CIT cannot address all issues shown in Figure 1, where LLM underperformance may arise from a lack of requisite knowledge rather than unfamiliarity with the task schema Ren et al. (2023); Zhang et al. (2023b). In contrast, CPT better addresses such knowledge deficits.
@@
However, using CPT to supplement the model with lacking knowledge still presents several challenges. First, collecting CPT data is challenging as identifying the lacking knowledge a priori while ensuring a sufficiently large data size is difficult. Blind data collection and insufficient data volume may introduce irrelevant or redundant information, weaken generalization, and increase overfitting risks Zhang et al. (2024); Gekhman et al. (2024b). Second, efficiently utilizing CPT data and mitigating degradation are difficult. As Jiang et al. (2024) note, even at minimal perplexity, not all knowledge can be extracted—a ‘perplexity curse’ that limits effective data utilization. Moreover, compared to supervised fine-tuning, the larger parameter shifts in CPT increase the risk of performance degradation.
@@
To address these challenges, common data collection strategies rely on the latest Wikipedia snapshots or new domain academic papers Cossu et al. (2022); Jin et al. (2022). However, these methods lack focus and efficiency and fail to dynamically update model deficiencies. Additionally, recent work has explored using highly relevant QA pairs for CIT before CPT Jiang et al. (2024) or converting raw CPT corpora into CIT instructions (e.g., reading comprehension instructions) Cheng et al. (2024b), thereby reducing degradation and enhancing knowledge absorption. Nevertheless, their dependence on advanced LLMs and the complexity of mining patterns limit their practical applicability.
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, which: (1) introduces a practical pipeline to collect CPT data derived from model mistakes, and (2) employs a novel training paradigm that integrates CIT and CPT to efficiently utilize these data. Together, these innovations enable continual model improvement while mitigating performance degradation.
@@
To enhance the focus and efficiency of data collection, we draw inspiration from the human learning skill of ‘summarizing mistakes’. We argue that a model’s mistakes in ‘exams’ reflect its inherent knowledge deficiencies. Accordingly, we collect background information from the internet related to these identified mistakes, directly addressing the specific knowledge deficiencies they reveal. To further optimize CPT data utilization and mitigate forgetting, we construct training sets integrating CIT and CPT data in parallel. Coupled with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al. (2022), CEM supports iterative, continual model evolution. Across extensive experiments, CEM yields substantial gains for multiple models, increasing accuracy by up to 29.63%.
@@ 2 Related Work
Continual pre-training enhances domain knowledge through two approaches: (1) domain-incremental pre-training, which accumulates knowledge across multiple domains, and (2) domain-specific continual learning, which refines a general model into a domain expert by training on domain-specific datasets and tasks. In domain-incremental pre-training, Cossu et al. (2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various asks. Ke et al. (2023) introduced a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge. Notably, there has been research on continual pre-training for specific domains such as finance Xie et al. (2023) and e-commerce Ma et al. (2023).
@@
Task-incremental CIT aims to continually fine-tune LLMs on new task-specific instructions. While a straightforward approach involves generating new instruction-tuning data and fine-tuning LLMs directly Wang et al. (2023), this method can lead to catastrophic forgetting of previously learned knowledge and skills Kotha et al. (2024).
@@ 3.1 CEM Method
(1) Search Engine: Search engines are known for high relevance matching and diverse content. We search for extracted knowledge points using Bing and select the top-k results based on relevance, scraping the main textual content as a supplemental corpus. To prevent LLMs from simply memorizing answers, we employ a 13-gram matching to filter out any results that contain both the original question and its answer. Moreover, we use newspaper3k [^3] https://github.com/codelucas/newspaper and readability [^4] https://github.com/buriy/python-readability to exclude irrelevant and spam content.
@@
It should be noted that supplemental training should not be endless, as it may lead to catastrophic forgetting and diminishing returns in accuracy improvement (i.e., ‘perplexity curse’). Subsequent experiments explore the impact of the number of training rounds on model performance.
@@ 4.2 Datasets
We conduct experiments on two question-answering datasets, Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024a). To evaluate CEM’s out-of-domain generalization, we apply CEM on Xiezhi as the in-domain dataset and use CMMLU and C-Eval Huang et al. (2023b) as out-of-domain benchmarks.
@@ 4.4 Metrics
ER: The Enhancement Rate measures the relative improvement brought by CEM on a specific task compared to the initial performance. It is calculated by $ER=\frac{A_{k}^{*}-A_{k}^{0}}{A_{k}^{0}},$ where $k$ denotes the task using CEM, $A^{0}_{k}$ is the initial accuracy, and $A^{*}_{k}$ is the accuracy after CEM.
AFR: The Average Forgetting Rate Wang et al. (2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance. It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{A_{i}^{0}-A_{i}^{*}}{A_{i}^{0}}\right)$, where $A^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $A_{i}^{*}$ is the accuracy of tasks $i$ after applying CEM on task $k$.
@@ 5.1 Main Results
Firstly, CEM leads to significant accuracy improvements across all models and tasks. All variations of CEM substantially enhance model performance in both the in-domain (ID) Xiezhi task and OOD tasks. For instance, the Llama3-8B-Instruct model achieves accuracy gains of up to 29.63% on the Xiezhi task, 10.92% on the CMMLU task, and 5.40% on the C-Eval task when using CEM-R. This highlights the effectiveness of the CEM method in addressing the model’s knowledge gaps and its capacity to generalize this knowledge to unseen QA tasks.
@@ 5.2 Data Source Analysis
We analyze the impact of the Supplementary Corpus from different sources—Wikipedia, Bing, and their concatenation—on CEM’s performance. To further validate CEM’s applicability, CMMLU is included as an ID task scenario with data proportions consistent with the main experiment.
@@
We find that the CEM method shows robust compatibility across different data sources, consistently improving model performance. However, integrating multiple sources does not necessarily outperform using a single source. As shown in Figure 3, no significant differences in CEM effectiveness emerge when employing Wikipedia, Bing, or their combination. Although the Mix dataset is larger, it often yields lower accuracy than Wiki or Bing, potentially due to data redundancies. Moreover, an excessively large supplemental corpus can dilute the influence of CIT data, reducing its capacity to mitigate catastrophic forgetting, enhance knowledge extraction, and prevent overfitting.
@@ 5.3 Mechanistic Analysis of Extractive and Review Instruction Components
To clarify the underlying mechanisms by which Extractive Instruction and Review Instruction affect model performance, we conduct experiments on CuteGPT-13B-ift and examine the R2W and W2R metrics. Additionally, we explore the effects of varying $\alpha$ (with values $[0,0.5,1]$) on CEM-R to determine whether more diverse and complex Review Instructions yield improved results. The corresponding results are illustrated in Figure 4 and Figure 5.
[…]
We summarize the key observations and insights derived from these experiments:
(1) CEM-E and CEM-R outperform their ablation groups with General Instruction, demonstrating the effectiveness and positive impact of Extractive and Review Instruction on model performance. As shown in Figure 5, the Extractive Instruction enhances W2R performance: compared to CEM-P, CEM-E achieves improvements of 1.45%, 3.22%, and 1.26% across the three models, indicating more effective error correction. In addition, the Review Instruction leads to a significant reduction in the R2W metric, indicating a reduced tendency for the model to forget previously acquired correct knowledge. For example, using CEM-R1, Llama3-8B-Instruct’s R2W performance decreases by 5.61% and 4.29%, respectively compared to CEM-P and CEM-E.
@@ 5.4 Multiple Iterations Analysis
To evaluate the iterability of the CEM method and its potential for catastrophic forgetting during multiple training rounds, we conduct three rounds of CEM training on CuteGPT-13B-ift for the Xiezhi task, using consistent training strategies and non-repetitive CIT data of equivalent size in each round. Specifically, the CEM method is applied to the Xiezhi task, while HotpotQA and GSM8K are used to assess reasoning and mathematical skills. The ER metric evaluates performance improvements on Xiezhi, and the AFR metric monitors performance degradation in other tasks. Initially, 2,000 instances from each dataset are sampled for pre-finetuning to establish a baseline. To address forgetting, we implement a random replay strategy, replacing General Instructions with 400 randomly replayed instances from HotpotQA and GSM8K in each training round.
@@
Figure 6 shows that even after three rounds of CEM, performance on Xiezhi improves significantly and maintains an upward trend. The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 75.0% in most cases, highlighting its iterative potential. However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking instructions introduced by random replay), but more rounds lead to greater forgetting, limiting further iteration, as seen when CEM-R reaches an ER of 62.4% in round three without replay. CEM-R, with its larger data capacity, achieves significant performance improvements but shows the greatest decline in other tasks, likely due to the increased volume of CPT data diluting the replay effect. Combining CEM with other advanced replay-based methods and data balancing warrants further investigation.
@@ 6 Conclusion
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLMs on specific tasks by identifying and addressing internal knowledge deficiencies revealed by models’ mistakes. CEM employs a pipeline that leverages model mistakes to gather extensive and targeted CPT data, combined with a novel parallel training set construction paradigm integrating both CIT and CPT data for continuous training. Experiments on various open-source LLMs show that CEM significantly enhances models’ answer accuracy and supplements their internal knowledge deficiencies. These improvements effectively generalize to out-of-domain tasks. Furthermore, when integrated with replay-based methods, CEM demonstrates substantial potential for sustained evolution across multiple iterations while mitigating performance degradation.